# Surface Reconstruction Using Rotation Systems

### v1

@@ 2.1. Euler Operators
where $g$ is the genus of the object represented by the mesh. This formula must be satisfied by any closed, manifold mesh. As such, it has been used as an invariant in solid modeling operations as discussed in e.g. (Mantyla and Sulonen, 1982). Typically, operators which maintain the Euler-Poincare formula invariant are called Euler operators. In the following, we will describe the two Euler operators we will need for our surface reconstruction algorithm and how the rotation system of the mesh is affected by these operations. The operators are illustrated in Figure 3.
@@ 2.1.1. Splitting a face by edge insertion
A face is split by creating an edge, $e$, between two vertices, $v$ and $w$. Assuming $v$ and $w$ belong to the same face, $f$, this operation splits $f$ into two faces. Thus, it is clear that (1) holds since $|E|$ and $|F|$ both increase by one.
@@ 2.1.2. Adding a handle by edge insertion
A topological handle is added by creating an edge, $e$, between a vertex $v$ that belongs to face $f_{1}$ and a vertex $w$ belonging to a different face, $f_{2}$. In this case, $f_{1}$ and $f_{2}$ are merged into a single face which we can see as a tube connecting the holes left by removing $f_{1}$ and $f_{2}$. In other words, this operation adds a handle to the mesh, increasing $g$ by one. (1) still holds since $|E|$ and $|F|$ increase and decrease by one, respectively.
@@
Figure 3. Before (top) and after (below) illustration of the edge insertion operation both in the case (left) where a face $f$ is split into two faces and in the case (right) where two faces, $f_{1}$ and $f_{2}$, are merged into a single face.
@@ 3. Method
Our reconstruction method initially forms a tree from the input points and then proceeds to triangulate the tree. The method can be described in terms of three steps, as shown in Figure 1.
  1. (1)
    Initially, we form a graph, $G$, from the input point cloud $V$, $V\subseteq\mathbb{R}^{3}$, by connecting each point to its $k$ nearest neighbors ($k$NN) within a given max radius, $r$. Based on $V$, we estimate the normal by applying principal component analysis (PCA) and correct the orientation through a normal-angle-weighted minimum spanning tree (MST). Afterward, we compute a rotation system (RS) that stores the cyclic ordering of the outgoing edges for each vertex. Based on the $k$NN graph, a distance-weighted MST, $T$, is produced from this graph $G$.
  2. (2)
    Next, to increase the connectivity of $T$, we insert edges from $G$ that connect leaf vertices to other vertices which are potentially far away in terms of graph distance but belong to the same face. Before making a connection, we perform both a topological check and checks of the geometric validity and quality of the connection.
  3. (3)
    Finally, we create the mesh $M$. $M$ is initialized to $T$ augmented with the additional edges. We split triangles from the faces of $M$ by greedily adding edges from $G$ until no more valid triangles can be formed.
In the following, we will describe each of these steps in more detail and then discuss how the algorithm can be generalized to allow for meshes of arbitrary genus.
@@ 3.1. Initialization
As stated earlier, the initial graph $G$ is derived from the input point cloud. Vertices are formed from the points, and each vertex is connected to its $k$ nearest neighbors ($k$ and other parameters are listed in Table 1). Based on $G$, two minimum spanning trees, a normal-angle-weighted MST and a distance-weighted MST, are generated for later use.
Unoriented normals are computed using the PCA method, but in order to construct the correct rotation system, we need the normal orientations. These we obtain from the normal-angle-weighted MST. Edges in this MST keep a weight calculated using the normals of involved vertices (Hoppe et al., 1992),
(2) $w_{u,v}=1-|\mathbf{n}_{u}^{T}\mathbf{n}_{v}|$
In other words, edges between vertices $u$ and $v$ with small angles between their respective normals have a small weight. Starting from an arbitrary vertex, a consistent normal orientation is propagated to all vertices along the edges of the normal-angle-weighted MST.
Figure 5. RS of an example neighborhood. Taking $v_{2}$ as an example, it is projected on the disk and then an angle $\alpha$ is calculated as the radian of vertex $v_{2}$.
The distance-weighted MST, $T$, is used as the initial structure of the mesh. It is generated based on the Euclidean distance as the edge weight. At the same time, we build the rotation system (RS) which determines the cyclic ordering of the neighbors of all vertices. As shown in Figure 5, every neighbor of $p_{r}$ is projected onto its tangent plane (based on the oriented normals). Starting from a given (but arbitrary) reference direction, the angles of neighbors are calculated, and a cyclic ordering is computed by sorting the neighbors according to these angles. Effectively, this produces the $\rho$ operator. The RS is computed on the original graph $G$, but not all edges are a part of $T$. Given an halfedge, $h\in T$, $\rho(h)$ yields the next halfedge in the cyclic ordering of outgoing halfedges belonging to $G\cap T$. As $T$ is updated by inserting edges from $G$, so is the $\rho$ operator.
@@ 3.2. Edge insertion
Before triangulating the MST $T$ to form a mesh, an edge insertion step is needed to increase the connectivity of $T$.
Figure 6. An example showing the necessity of edge insertion. The left figure presents the position of input 2D points. The reconstruction result in the middle is generated without edge insertion, where the blue lines are the MST. The figure on the right depicts an ideal reconstruction result. This is achieved by the edge insertion before triangulation, the red edge in this plot.
While the MST is a planar graph by virtue of being a tree, directly proceeding to triangulation generally leads to a poor result where ill-shaped and self-intersecting triangles are formed. Figure 6 illustrates why this can happen. A 3D illustration on Stanford Armadillo is shown in Figure 7. The lack of connectivity of $T$ leads to overlapping geometry in this example. Compared with the triangulation stage, edge insertion is more time-consuming with both topology and geometry tests. Hence, establishing possible connections on all leaf nodes emerges as a viable strategy, ensuring ample connectivity within a reasonable execution timeframe.
Figure 7. A 3D example of triangulation without edge insertion, where the blue edges are the MST. A random origin of the crack is zoomed in, which is caused by the limited connectivity of $T$.
For every leaf node, we take its $k/2$ nearest neighbors as connection candidates. The closest candidate passing specific tests will be connected.
Two mandatory tests are employed on all leaf nodes considered as end-points for new edges. The first is a topological test, which ensures that the edge insertion does not change the topology of $T$ that should remain planar (although we relax this constraint later). The second is a geometry test, which ensures that the edge insertion does not lead to self-intersecting triangles.
In practice, we also use a quality test to avoid geometrically poor configurations. Also, the quality check increases efficiency by rejecting invalid edges via more lightweight checks. More details will be introduced in Section  4.2.1.
@@ 3.2.1. Topology Test
As mentioned, a face in a graph is defined as an orbit, i.e. the set of edges obtained by repeated application of $\tau$. The topology of the graph will not change as long as inserted edges only split faces rather than connecting different faces. Examples can be found in Sec. 5.1. From a practical point of view, we store an identifier of the associated face with each halfedge. These identifiers are stored in a tree structure for efficiency concerns (for more information, please refer to Sec. 4.3). For a candidate edge, $\{u,v\}$ we check both that $u$ and $v$ belong to the same face, $f$, and also that when $(u,v)$ is inserted in the cyclic ordering of outgoing edges at $u$ and $(v,u)$ in the corresponding ordering at $v$, the corners that are split also belong to $f$ as mentioned in Sec. 2.1.1.
@@ 3.2.2. Geometry Test
The geometry test inspects the local geometry around a given vertex within the tangent plane to avoid self-intersecting or overlapping triangles. As illustrated in Figure 8, two point sets, a query set and a rejection set, are retrieved based on Euclidean distance. Each existing edge in the rejection set, such as edge $e$ in the figure, will be used to reject candidate connections if intersections occur. The radius of the rejection set should be large enough to include all possible rejection edges. Thus, it is calculated with (3), where $r_{\mathrm{qry}}$ stands for the radius of the query set while $L_{\mathrm{max}}$ represents the maximal existing edge length.
Figure 8. A geometry test applied to vertex v. The green points are the query set and the blue points together with the green points are the rejection set. The red dashed line represents the candidate edge insertion and the black edge e is an existing edge. In the left case, the candidate end vertex u will be rejected because of the existence of e. The right configuration shows the extreme situation where the radius of the rejection set reaches the maximum.
(3) $r_{\mathrm{rej}}=r_{\mathrm{qry}}+L_{\mathrm{max}}$
When extending this check to 3D, an additional requirement about vertex normal orientation on the query/rejection set selection is applied, to deal with thin structures or areas with high curvature. If a vertex $u$ and root $v$, with normal $\mathbf{n}_{u}$ and $\mathbf{n}_{v}$, have a negative inner product on normal, i.e. $\mathbf{n}_{u}\cdot\mathbf{n}_{v}<0$, then vertex $u$ is not included in either point set.

### v2

@@ 2.2. Euler Operators
where $g$ is the genus of the object represented by the mesh, and the value $2(1-g)$ is known as the Euler characteristic. This formula must be satisfied by any closed, manifold mesh. As such, it has been used as an invariant in solid modeling operations as discussed by e.g. Mantyla and Sulonen (1982). Typically, operators that maintain the Euler-Poincare formula invariant are called Euler operators. In the following, we will describe the two Euler operators we will need for our surface reconstruction algorithm and how the rotation system of the mesh is affected by these operations. The operators are illustrated in Figure 4.
@@ 2.2.1. Splitting a face by edge insertion
A face is split by creating an edge, $e$, between two non-adjacent vertices, $v$ and $w$. Assuming $v$ and $w$ belong to the same face, $f$, this operation splits $f$ into two faces. Thus, it is clear that (1) holds since $|E|$ and $|F|$ both increase by one.
@@
Figure 4. Before (top) and after (below) illustration of the edge insertion operation both in the case (left) where a face $f$ is split into two faces and in the case (right) where two faces, $f_{1}$ and $f_{2}$, are merged into a single face $f$. Note that the new face, $f$, is a tube that connects the boundaries of $f_{1}$ and $f_{2}$ which is hard to convey adequately in a 2D drawing.
@@ 2.2.2. Adding a handle by edge insertion
A topological handle is added by creating an edge, $e$, between a vertex $v$ that belongs to face $f_{1}$ and a vertex $w$ belonging to a different face, $f_{2}$. In this case, $f_{1}$ and $f_{2}$ are merged into a single face which we can see as a tube connecting the holes left by removing $f_{1}$ and $f_{2}$. Intuitively, one can think of the new face as a piece of paper that has been rolled into a cylinder: the boundary curves of the cylinder correspond to the boundaries of $f_{1}$ and $f_{2}$ whereas $e$ corresponds to the edge where the opposite sides of the paper meet.
Topologically, this operation adds a handle to the mesh, increasing $g$ by one. Equation (1) still holds since $|E|$ and $|F|$ increase and decrease by one, respectively.
@@
Figure 5. $RS$ of an example neighborhood. Taking $v_{2}$ as an example, it is projected on the disk. Then an angle $\alpha_{2}$ is calculated as the radian of vertex $v_{2}$.
@@ 3. Method
Figure 6. Pipeline overview on synthetic letter point clouds. High-genus reconstruction follows the bottom pipeline, where $g$ stands for the genus number. Given the prior knowledge of a genus-0 shape, the shorter pipeline depicted at the top is employed. In the pipeline shown at the bottom, the red edges in the third inset denote the newly added handles. Concurrently, the path formed by the blue edges represents the shortest path between the two end vertices before the red edge is introduced.
The essence of our algorithm is to construct a tree from the input points and build a triangle mesh from this starting point. As mentioned, a tree is known to be genus-0, and we ensure that this property is maintained while regular edges are added. As an optional part of the algorithm, handle edges can be added to increase the genus of the reconstructed object.
In the following, we will discuss all stages in greater detail and we refer the reader to Figure 6 and Algorithm 1 for an overview of our pipeline.
Input : pts, $r,k,\theta,n$, normals
pts_proj $\leftarrow$ pts
if points are noisy then
       pts_proj $\leftarrow$ project(pts, normals) // c.f. Section 3.4
end if
$\Gamma\leftarrow$ build_knn_graph(pts_proj, $r$, $k$)
for $G$ in components($\Gamma$do
       $CO\leftarrow$ build_circular_ordering($G$, normals)
       MST $\leftarrow$ init_MST($G$)
       $(M,RS)\leftarrow$ (MST, build_RS(MST, $CO$))
       $Q\leftarrow$ sort_edges($G$)
       for edge in $Q$ do
             if $\mathrm{topology\_test}(M,RS,\mathrm{edge})$ then
                   if $\mathrm{geometry\_test}(M,RS,\mathrm{edge},\theta,$ pts_proj) then
                         $\mathrm{insert\_edge}(M,\mathrm{edge})$
                         $RS\leftarrow\mathrm{update\_RS}(RS,CO$, edge)
                   end if
             end if
       end for
      if handle connection enabled then
             $\mathrm{connect\_handle}(M,RS,n)$
             $\mathrm{triangulate}(M,RS)$
             // c.f. Section 3.3
       end if
      output_mesh($M,RS$, pts)
end for
Algorithm 1 Pseudocode for the main loop
@@ 3.1. Initialization
Given the input point cloud, we build a k-nearest neighbor ($k$NN) graph, $\Gamma$, where $k$ is a user-defined parameter and the Euclidean distance is used to define the neighborhood. Notably, since the connections are symmetrized for both incident vertices of an edge, a vertex may have more than $k$ neighbors. However, edges are also removed if the connected vertices $u$ and $v$ have normal vectors which point in significantly different directions, i.e. $\mathbf{N}_{u}\cdot\mathbf{N}_{v}<cos(\theta)$, where $\theta$ is also a user-defined parameter. This test helps us avoid connecting vertices on either side of thin structures. Finally, edges longer than $r$ times the average edge length are removed as outliers. We refer the reader to Section 4.5 for further discussion of $r$ and the two other parameters.
For input point clouds that are known to be free of noise, we can use $\Gamma$ directly in the reconstruction process described below. However, for noisy inputs, we compute a filtered point set and a corresponding graph as described in Section 3.4.
The next step is to compute and iterate through all the connected components of $\Gamma$ and reconstruct a triangle mesh for each component, $G=(V,E_{G})$.
Figure 7. Topology check for the candidate edge $\left\{u,v\right\}$. In this example, the candidate edge splits $f_{1}$ within the right corner. Thus, it passes the topology check.
Based on the vertex normals, we now establish a cyclic ordering, $CO$, of the edges incident on all vertices of $G$. The construction of this ordering for a vertex $u$ with neighbors, $v_{i}$, where $\left\{u,v_{i}\right\}\in E_{G}$ is shown in Figure 5. Each neighboring vertex, $v_{i}$, is projected into the plane defined by the position $\mathbf{p}_{u}$ and the normal, $\mathbf{N}_{u}$, of vertex $u$. We compute the angle of each projected vertex relative to an arbitrary reference direction and thus obtain a cyclic ordering of the neighbors of $u$. While our approach thus relies on vertex normals, the normal computation is not an intrinsic part of the scheme, and our code allows for user-provided normals. In both respects, our approach is similar to many other reconstruction methods.
As the next step, a graph, $M=(V,E_{M})$, is initialized to the minimum spanning tree (MST) of $G$, and we initialize the rotation system, $RS$, such that the edges $E_{M}\subset E_{G}$ have the same cyclic order in $RS$ as they do in $CO$. Only if this is true do we consider the rotation system to be consistent with the input normals.
At this point, $M$ defines the vertices and edges of a mesh, and $RS$ defines the faces in terms of the operator $\rho$. Since $M$ is a tree initially, it consists of only a single face that can be traversed starting from an arbitrary halfedge using repeated application of $\tau=\rho\circ\iota$. In the next phase, additional edges from $E_{G}$ will be added to $E_{M}$. As $E_{M}$ expands, $RS$ will be updated while ensuring that it remains consistent as defined above.
In addition to the rotation system, we need to be able to keep track of the face associated with each halfedge. A balanced binary tree (not to be confused with the minimum spanning tree of $G$) is used for this purpose. From the tree, the face associated with each halfedge can be retrieved, and we can update the tree efficiently when a new edge is inserted, splitting one of the faces into two faces (for more information, please refer to Section 4.1).
As the final initialization step, all edges in $G$ are sorted in order of increasing length and stored in the list $Q$ since we will consider edges in this order.
@@ 3.2. Edge Insertion
In the edge insertion stage, we iteratively pick edges from $Q$ and check whether they are valid. Valid edges are inserted into $E_{M}$, thereby splitting a face of $M$ as illustrated in Figure 4. An edge is considered valid if it passes two tests, the topology test and the geometry test, which, respectively, ensure that $M$ remains a genus-0 graph and locally a pp graph. In other words, the first test is required to ensure that we do not make any topological errors and the second test ensures we avoid flipped triangles and overlapping edges in the mesh.
@@ 3.2.1. Topology Test
This test first ensures that the endpoints of the candidate edge belong to the same face, $f$. Next, it is verified that both of the corners which are split when the candidate edge is inserted belong to $f$. We refer the reader to Section 2.1 for the definition of corner.
In order to carry out the topology test in practice, we store the face associated with each halfedge. When a candidate edge $\left\{u,v\right\}$ is considered, we check that all the halfedges before and after $\{u,v\}$ in the rotation system, i.e. $h_{1},h_{2},h_{3}$ and $h_{4}$, are associated with the same face, $f_{1}$ as shown in Figure 7.
Figure 8. Example for the necessity of topology test. The left mesh is reconstructed with both tests, while the right one is reconstructed without the topology test. A red triangle is formed which results in an undesirable hole.
It is clear that $M$ remains genus-0 if the edge candidate passes the topology test because $|E|$ and $|F|$ both increase by one. Consequently, the left-hand side of the Euler-Poincare formula (1) does not change. This also implies that the topology test ensures that handles are not inserted at this stage, which is why our method enables reconstruction with topology control.
@@ 3.2.2. Geometry Test
Figure 9. Geometry test for edge {$u,v$}. With a search ball of radius $\frac{1}{2}l+L_{\max}$ and centered at the midpoint of the edge candidate, this test will check if any existing edge associated with vertices inside this search ball intersects with the candidate edge within the tangent plane.
The geometry test ensures that a 2D projection of a local neighborhood of the graph remains a plane graph, i.e. a pp graph, when a new edge is added. As depicted in Figure 9, the local neighborhood is defined as a ball with a radius of $\frac{1}{2}l+L_{\max}$, where $l$ is the length of the edge candidate and $L_{\max}$ is the maximum length of any edge in $G$. The plane, $P$, passes through the midpoint of the candidate edge and its normal is defined as the average of the normals at the endpoint vertices.
Any edge for which it is true that either endpoint belongs to the local neighborhood is projected into $P$ and we test for intersection with the candidate edge. If there is an intersection, we reject the candidate edge.
We stress that both the topology test and the geometry test are required. Even if topology control is not desired, the topology test is needed in cases where the tangent plane changes rapidly. In such cases, the geometry test may pass for edges that connect different faces as illustrated in Figure 8.
Figure 10. Example for the necessity of geometry test. The left mesh is reconstructed with both tests, while the right one is reconstructed without the geometry test. The graph remains genus-0, but no edge candidate left can pass the corner check of the topology test. Thus, a hole is left here and the same reason for any other place of the reconstructed mesh.
Likewise, the topology test is not sufficient in itself. As shown in Figure 10, there are cases where an edge that passes the topology test would lead to undesirable configurations, such as intersections, which do not increase the genus number. Fortunately, these cases are easily dealt with by the geometry test.
Candidate edges that pass both the geometry and topology test can safely be inserted. However, the two tests do not guarantee the quality of the triangles. In order to prevent skinny triangles, i.e. caps or needles with one angle close to $180^{\circ}$ or $0^{\circ}$, respectively, we also perform a quality check before insertion. This is discussed in Section 4.2.
It bears mentioning that both the topology test and the geometry test depend on the normal vectors. In the case of the topology test this means that very noisy normal vectors could result in fewer edges that pass the test, but it does not result in edges erroneously passing the test. In the case of the geometry test, no strong guarantees are provided, but normal vector noise would generally also lead to edges failing this test. Ultimately, this means that the deleterious effects of noisy normal vectors can be summed up as being fewer connections which, for significant levels of noise, can lead to holes in the mesh. We investigate the effect of synthetic noise on the normal vectors in Section 5.5.2.
3.3. Handle Connection & Triangulation
Up to this point, $M$ is guaranteed to be of sphere topology, but some of the faces of $M$ likely consist of many edges. Given such a large face, $f_{1}$, and a vertex, $v$, that belongs to $f_{1}$. $v$ may either lack connections to other vertices of $f_{1}$ because there is a gap in the input point cloud or because the unused edges of $G$ which are incident on $v$ connect it to a different face, $f_{2}$. In the former case, we would want to leave $f_{1}$ as a hole, but in the latter case, we may want to insert a handle edge between $f_{1}$ and $f_{2}$
To merge $f_{1}$ and $f_{2}$, we need to close the cracks. A crack is constituted by a pair of faces $f_{1}$ and $f_{2}$ such that halfedges on one side are associated with $f_{1}$, while halfedges on the other side are associated with $f_{2}$. We create a set of handle edges by finding vertices belonging to a face $f_{1}$ that are connected to a different face and whose corner (i.e. the corner that would be split if the edge were inserted) has an angle greater than $\pi$. This restriction on corner angle aims to avoid skinny triangles as a quality concern. A summary of quality concerns is provided in Section 4.2 Such handle edges are now inserted if they pass the geometry test described above. The topology test does not apply since it would trivially fail for handles. On the other hand, small, undesired handle candidates are often present locally – in particular on noisy meshes. Thus, we need a different test for handle insertion, and the following heuristic test is employed.
When a handle edge is considered for insertion, we can find the shortest path between the endpoints that consists of edges already in $M$. If the handle is well-sampled, this path consists of a significant number of edges. For instance, consider the letter A in the lower row of Figure 6: it is clear that the path connecting the endpoints of the handle edge seen in the inset must go around the hole in the A. Thus, we use a threshold, $n$, on the number of edges in the shortest path between the incident vertices of the handle candidate edge, to cull undesirable handles. A higher value of $n$ corresponds to fewer handles being connected. For the input point clouds we tested, an effective setting for $n$ is $50$.
After connecting all handle edges that pass the above test, a final triangulation is performed. At this point, most of the faces are indeed triangles, but since we reject some edges due to quality concerns both during edge insertion and handle insertion (c.f. Section 4.2), not all faces are fully triangulated. After handle insertion, our face data structure is no longer valid since it can only be efficiently updated when faces are split and not merged. Fortunately, the rotation system is unaffected, and the triangulation only performs ear clipping, i.e. it connects two vertices known to belong to the same face. More details are provided in Section 4.3.
3.4. Reconstruction from Noisy Point Clouds
Scanned point clouds are inherently noisy since scanning is a physical measurement process. Moreover, several sub-scans are often combined during reconstruction. These may not be perfectly aligned, introducing another source of high-frequency noise.
Noise can be both tangential and normal to the surface. While tangential noise does not affect the reconstruction, normal noise can result in large Euclidean distances between points that are close in the tangent plane. To mitigate this issue, two measures are taken against noisy data.
First, we use the provided normals to project all points onto their tangent planes and compute the projected distances which are then used instead of Euclidean distance. To define the projection distance, we need a preliminary definition: for an edge $e=\{u,v\}$, $e_{//v}$ refers to the projection of $e$ on the tangent plane of $v$. The same applies to $u$. The projection distance $|e|_{p}$ is now,
(2) $|e|_{p}=(|e_{//v}|+|e_{//u}|)/2\enspace.$
However, the kd-tree used for computing the initial graph $G$ is, of course, unaffected by the tangential distances and still based on the Euclidean distance. This can lead to both spurious and missing connections. Therefore, we also project the input points to the average tangent plane of their k-neighborhood and rebuild the kd-tree. Inspired by (Digne et al., 2011), we restore the input points to their original position before the mesh is output.
It is reasonable to ask why we do not simply compute the Euclidean distance after projecting the points. The reason is that even after the above smoothing, two connected points never lie exactly in each others tangent plane. Hence, there is a meaningful difference between the Euclidean distances and the tangent plane projected distance, and we obtain the best results by using the pure tangent plane distance.
4. Implementation
In this section, we discuss the practical aspects of the method. These aspects include some details pertaining to the binary tree data structure used to store faces, some quality concerns, details related to triangulation after handle connection, the time complexity of each stage, and the parameters used in the method.
4.1. Face identifier maintenance
In order to efficiently perform the topology test described in Section 3.2.1, we need a data structure that allows us to quickly determine if two halfedges lie on the boundary of the same face, while also supporting fast updates as edges are inserted and faces split.
Using the Euler Tour technique (Henzinger and King, 1999), we store the sequence of halfedges of each face in a balanced binary tree with implicit keys. To query if two half-edges lie on the same face, we can simply traverse up to the roots of the associated binary trees and compare, allowing for $O(\log|V|)$ queries.
Storing the faces in this manner also allows for the use of standard binary tree operations, such as splits and joins. By rebalancing while performing these operations, they can be executed in $O(\log|V|)$ time while maintaining balance. When inserting a new edge, we perform at most two splits to obtain binary trees representing the sequences of halfedges that bound the newly split face, and at most three joins to incorporate the new half-edges in these sequences. Thus, we can maintain the faces under the insertion of new edges in $O(\log|V|)$ time as well.
4.2. Quality concern
To ensure the quality of the output mesh, we impose some constraints throughout the reconstruction procedure.
First, if an edge candidate results in a triangle with an angle larger than $175^{\circ}$ or smaller than $5^{\circ}$, we reject the candidate. Often, better candidate edges are connected later, and if this does not happen, the final triangulation guarantees that the output is a triangle mesh.
When the reconstructed objects are of genus $>0$, we take an additional measure. The issue is that handle edges, which connect different faces, are not inserted during edge insertion. This, in turn, means that long edges that connect two vertices on the same side of a crack (please refer to Figure 6 for an illustration of a crack) may be selected. However, this could again lead to skinny triangles. Thus, for a shape that is not genus 0, we are more conservative, selecting just the first $\frac{2}{3}$ of the sorted edges from $G$ in the edge insertion stage, leaving the remaining edge insertion to the triangulation stage.
Finally, during handle connection, we only consider handle edges that connect to vertices at a corner with an angle greater than $\pi$. Again, this is to avoid skinny triangles.